a fruit vendor has 990 apples and 945 oranges. He packs them into basket.each basket contains only one of the two fruits but in equal number. Find the number of fruits put in each basket in order to have minimum no. of baskets?
Answer by ewatrrr(24785)

990 = 22*45
945 = 21*45
45, the number of fruits put in each basket in order to have minimum no. of baskets
(45 the lowest common denominator for 990 and 945)
990 = 9*5*22
945 = 9*5*21